# Better Safe Than Sorry: Enhancing Arbitration Graphs for Safe and Robust Autonomous Decision-Making

### v2

@@ Abstract
0.82(0.09,0.94) This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I-B State of the art
Arbitration graphs, combining subsumption and object-oriented programming, enhance reusability and system clarity through modularity and functional decomposition. Used in robotic soccer [lauerCognitiveConceptsAutonomous2010] and automated driving [orzechowskiDecisionMakingAutomatedVehicles2020], these graphs employ behavior components to interpret situations and plan actions, while arbitrators select the most suitable behaviors.
@@ II-A Decision-Making
Decision-making is crucial in the planning module of a robot, determining commands based on the current situation. Methods range from graph-based techniques like A*, PRM*, and RRT* to probabilistic and machine learning methods. This work, however, focuses on rule-based methods, such as FSMs, BTs, or arbitration graphs, which address decision-making through discrete state or mode transitions. Refer to [schwartingPlanningDecisionMakingAutonomous2018, yurtseverSurveyAutonomousDriving2020, gammellAsymptoticallyOptimalSamplingBased2021] for a detailed overview of available approaches.
FSMs, first used in hardware design and theoretical computer science, represent behavior modes with transitions triggered by events [wagnerModelingSoftwareFinite2006]. Despite their simplicity, FSMs scale poorly and are difficult to modify due to the amount of transitions increasing exponentially with the number of states.
BTs, initially designed for game development [iovinoProgrammingEffortRequired2022], have been increasingly used in robotics since 2012 [bagnellIntegratedSystemAutonomous2012, ogrenIncreasingModularityUAV2012]. They separate behavior decision-making from execution using a tree structure [colledanchiseBehaviorTreesRobotics2018]. Internal nodes determine selection mechanisms, while leaves describe behaviors and conditions. Evaluated at a fixed frequency, nodes return their status as running, completed, or failed. Control flow nodes decide on further evaluations. Condition nodes check if their conditions are met without affecting the environment, while action nodes execute behaviors and return their status.
BTs generalize many architectures, such as hierarchical FSMs and decision trees [colledanchiseHowBehaviorTrees2017] excelling in modularity, hierarchical organization, reusability, responsiveness, and interpretability [colledanchiseBehaviorTreesRobotics2018]. Their flexibility allows reuse of individual behaviors [bagnellIntegratedSystemAutonomous2012]. The selection mechanism is intuitive and easy to follow during operation. However, extensive preconditions can make representations cumbersome, and safety as well as reliability depend significantly on node arrangement. These drawbacks are addressed by arbitration graphs.
Arbitration graphs originated in the context of robot soccer [lauerCognitiveConceptsAutonomous2010], integrating ideas from Brooks’ behavior-based subsumption [brooksRobustLayeredControl1986], knowledge-based architectures like Belief-Desire-Intention (BDI) [raoAbstractArchitectureRational1992], and programming paradigms such as object-oriented programming [stefikObjectOrientedProgrammingThemes1985].
@@ II-B Fault-tolerant/Robust Systems
Reliable systems aim to maintain performance despite potential errors, using metrics like error probability, mean lifespan, failure rate, and availability [echtleFehlertoleranzverfahren1990]. Terminology varies, but disturbances (“faults”) can lead to errors, potentially causing system failures.
Reliability measures include error prevention, removal, tolerance, and prediction [dubrovaFaultTolerantDesign2013]. Prevention and removal focus on design and development, while tolerance involves detecting and preventing operational errors. Prediction estimates future failures. Error tolerance involves diagnosing and handling errors—restoring faulty components, computing correct results despite faults, or removing faulty components [echtleFehlertoleranzverfahren1990].
@@ II-C Pac-Man
Pac-Man, the 1980 arcade game, involves navigating a maze to eat dots while avoiding ghosts. Depicted in Fig. 1 is an open-source implementation111github.com/indianakernick/EnTT-Pacman of the game that is being used to demonstrate the proposed method. We split the behavior of Pac-Man into the following behavioral components:
@@
The corresponding arbitration graph is shown in Fig. 2.
@@ III-A Detect Unsafe & Unreliable Behaviors
Algorithm 1 shows the generic arbitration algorithm from [lauerCognitiveConceptsAutonomous2010] that we have extended with a verification logic. The verifier is domain-specific and may run various checks on the behavior command. In Pac-Man for example, the verifier could check whether the command would lead to a collision with walls or ghosts. In autonomous driving, one could verify if the behavior output is free of collisions and respects traffic rules. Similarly, a robotic manipulator could verify if the command is within the workspace of the robot and does not exceed joint limits. More generally, the verifier could check if the format of the command is correct and respects the specification of the system. Integrating the verification step into the arbitration algorithm allows for the use of different verifiers at various levels within the arbitration graph. This enables a more fine-grained and adaptable safety concept. For example, verifiers can be customized for specific scenarios, or computationally intensive verifiers can be reserved for higher levels of the arbitration graph.
The verification step is embedded as follows: First, in Algorithm 1, the current situation is determined. Based on this, the root arbitrator determines its best applicable and safe action using the BestOption() function. For this, it determines the set of options which are applicable in the present situation . Like in the original arbitration algorithm, the options of an arbitrator can themselves be arbitrators or behavior components. The applicable options are then sorted into a descending list according to the underlying strategy. Now, for each option , it is checked whether its command withstands a verification . If so, this option is returned as the best applicable and safe option. If none of the options passes the verification step, the arbitrator returns the error value NO_SAFE_OPTION. Otherwise, the command returned by the root arbitrator can be executed and considered safe given the assumptions of the verifier .
@@ III-C Example: Safe Pac-Man
While the original arbitration graph shown in Fig. 2 might have been sufficient under normal conditions, it might fail in practice if a behavior component returns an unsafe or unreliable command. Consider the scenario depicted in Fig. 3. Since the ghosts are relatively far away and there is only one dot cluster left, the only applicable behavior is EatClosestDot. Ideally, this behavior has the intended effect and Pac-Man should move towards the last remaining dots. In our hypothetical scenario, however, a bug in the underlying path planning algorithm leads to erroneous results. Consequently, the behavior component fails to generate a valid command, causing the system to behave unpredictably or even crash.
Fig. 4 shows the arbitration graph from Fig. 2 extended with fallback layers. With EatClosestDot not returning a valid command, the system falls back to the newly added MoveRandomly behavior component. Moving Pac-Man in a random direction is a simple action which might help the primary behavior components to escape a deadlock and find a new valid command. In our scenario, however, MoveRandomly returns a command which would lead to a collision with a wall. Therefore, this command is rejected by the verifier as well, and the system falls back to the last resort behavior StayInPlace. While it might be impossible to complete the level with this behavior, the system remains in a predictable and safe state giving the primary behavior components a chance to recover.
@@ IV Safe and Reliable Behavior Arbitration for Automated Vehicles
Arbitration graphs can be applied to a wide range of real-world scenarios. In [orzechowskiDecisionMakingAutomatedVehicles2020], they have been used in the context of automated driving. There, a behavior component represents driving maneuvers of approximately to seconds such as lane changes, parking maneuvers, or crossing intersections.
@@
This section summarizes our work in [orzechowskiVerhaltensentscheidungFuerAutomatisierte2023], which extends [orzechowskiDecisionMakingAutomatedVehicles2020] by introducing a safety concept for arbitration graphs in the context of automated driving.
@@ IV-A Arbitration Graph — What to do?
Fig. 5 depicts a simplified arbitration graph for automated driving that will be used to showcase our safety concept. It contains three behavior components for basic lateral behaviors FollowLane, ChangeLaneLeft and ChangeLaneRight, a parking behavior ParkNearGoal and a stopping behavior EmergencyStop. The arbitration graph for a real-world application would include more behavior components and arbitrators in order to cover a wide range of driving maneuvers and scenarios. Refer to [waymo_safety_report_2020] for a detailed discussion about behavior competencies and operational design domains.
@@ IV-B Behavior Components — How to do it?
In our implementation, each behavior component has access to an environment model (situation ) and returns a trajectory (command ). The environment model contains information such as the ego vehicle state, other traffic participants, and the planned route. Fig. 6 illustrates an example scenario with the ego vehicle in blue, another vehicle in green, and the planned ego trajectory in dashed blue.
@@
See [orzechowskiVerhaltensentscheidungFuerAutomatisierte2023] for more details and other behavior components for automated driving.
@@ IV-C Verification — What is considered safe?
In order to achieve vehicle safety, we provide worst-case occupancy predictions of other traffic participants and extend the behavior command with a fail-safe trajectory [althoffSetBasedPredictionTraffic2016]. The fail-safe trajectory is designed to be safe under worst-case assumptions and can be used in fallback layers in case of a failure or too risky situation. As a result, each behavior component is responsible for generating both a planned trajectory and a fail-safe trajectory (see Fig. 6). The safety verifier checks if the occupancy of the fail-safe trajectory does not overlap with the worst-case occupancies of other traffic participants, i.e., the fail-safe trajectory is collision-free under worst-case assumptions.
@@ IV-D Fallback Layers — How to always drive safe?
In order to mitigate unsafe or unreliable behaviors, we extended the arbitration graph in Fig. 5 by the proposed verifiers and fallback layers highlighted by underlined text.
@@ IV-E Experiments
We validate the proposed concept in the automated driving simulator CoInCar-Sim [naumannCoInCarSimOpenSourceSimulation2018] using handcrafted but realistic driving scenarios from our real-world test track in Karlsruhe, Germany. We analyze two different use cases in [orzechowskiVerhaltensentscheidungFuerAutomatisierte2023]:
@@
Here, we focus on the latter, more critical case. Traditional decision-making approaches select maneuver options solely based on their preconditions. As a result, they might select unsafe maneuver options, if the preconditions have been designed too optimistically. Fig. 7 shows such a scenario, where the ego vehicle wants to change lanes, but another vehicle follows too closely in the target lane. We compare the evolution of the scene with and without verification in the decision-making process.
@@ IV-F Results
To provoke a risky situation, the invocation condition of ChangeLaneLeft was made overly optimistic. As a result, the arbitration graph without verification selects ChangeLaneLeft at leading to a collision at (Fig. 7).
With verification, the safety verifier detects the risk of a collision based on the worst-case occupancies of the other vehicle (Fig. 8). Consequently, ChangeLaneLeft fails the verification and the UrbanDriving arbitrator selects FollowLane as its next best option (Fig. 9). This makes the ego vehicle slow down and not change the lane until the green vehicle has passed (Fig. 10). A collision has been prevented while maintaining a smooth driving behavior.
@@ V Conclusion
This paper presented an extension to the arbitration graph framework that focuses on improving the safety and robustness of autonomous systems operating in complex, dynamic environments. It builds upon the strengths of arbitration graphs, which provide a flexible, scalable, and transparent decision-making framework for autonomous systems. By embedding a verification step into the arbitrators and adding structured fallback layers to the arbitration graph, the proposed method ensures that only verified and safe commands are executed.
The introduced method was demonstrated using a Pac-Man simulation, where the arbitration graph successfully maintained safe operation of the Pac-Man agent even in the presence of unexpected faults or bugs. Further validation was conducted in the context of autonomous driving, where the method demonstrated a reduction in accident risk and an improvement in system safety. These results underline the applicability of the proposed approach to real-world problems, confirming that arbitration graphs, when equipped with safety mechanisms, can effectively manage the complexities and uncertainties of autonomous decision-making.
@@
By explicitly defining the conditions under which a behavior component is considered safe, the responsibility for system safety is shifted to the verifiers used by the algorithm. This is a crucial advancement towards safe autonomous systems as the overall safety of the system now mainly depends on the assumptions made by these verifiers.

### v3

@@ I-B State of the art
Arbitration graphs, combining subsumption and object-oriented programming, enhance reusability and system clarity through modularity and functional decomposition. Used in robotic soccer [undef] and automated driving [undefa], these graphs employ behavior components to interpret situations and plan actions, while arbitrators select the most suitable behaviors.
@@ II-A Decision-Making
Decision-making is crucial in the planning module of a robot, determining commands based on the current situation. Methods range from graph-based techniques like A*, PRM*, and RRT* to probabilistic and machine learning methods. This work, however, focuses on rule-based methods, such as FSMs, BTs, or arbitration graphs, which address decision-making through discrete state or mode transitions. Refer to [undefb, undefc, undefd] for a detailed overview of available approaches.
FSMs, first used in hardware design and theoretical computer science, represent behavior modes with transitions triggered by events [undefe]. Despite their simplicity, FSMs scale poorly and are difficult to modify due to the amount of transitions increasing exponentially with the number of states.
BTs, initially designed for game development [undeff], have been increasingly used in robotics since 2012 [undefg, undefh]. They separate behavior decision-making from execution using a tree structure [undefi]. Internal nodes determine selection mechanisms, while leaves describe behaviors and conditions. Evaluated at a fixed frequency, nodes return their status as running, completed, or failed. Control flow nodes decide on further evaluations. Condition nodes check if their conditions are met without affecting the environment, while action nodes execute behaviors and return their status.
BTs generalize many architectures, such as hierarchical FSMs and decision trees [undefj] excelling in modularity, hierarchical organization, reusability, responsiveness, and interpretability [undefi]. Their flexibility allows reuse of individual behaviors [undefg]. The selection mechanism is intuitive and easy to follow during operation. However, extensive preconditions can make representations cumbersome, and safety as well as reliability depend significantly on node arrangement. These drawbacks are addressed by arbitration graphs.
Arbitration graphs originated in the context of robot soccer [undef], integrating ideas from Brooks’ behavior-based subsumption [undefk], knowledge-based architectures like Belief-Desire-Intention (BDI) [undefl], and programming paradigms such as object-oriented programming [undefm].
@@ II-B Fault-tolerant/Robust Systems
Reliable systems aim to maintain performance despite potential errors, using metrics like error probability, mean lifespan, failure rate, and availability [undefn]. Terminology varies, but disturbances (“faults”) can lead to errors, potentially causing system failures.
Reliability measures include error prevention, removal, tolerance, and prediction [undefo]. Prevention and removal focus on design and development, while tolerance involves detecting and preventing operational errors. Prediction estimates future failures. Error tolerance involves diagnosing and handling errors—restoring faulty components, computing correct results despite faults, or removing faulty components [undefn].
@@ II-C Pac-Man
Pac-Man, the 1980 arcade game, involves navigating a maze to eat dots while avoiding ghosts. Depicted in Fig.˜1 is an open-source implementation111github.com/indianakernick/EnTT-Pacman of the game that is being used to demonstrate the proposed method. We split the behavior of Pac-Man into the following behavioral components:
@@
The corresponding arbitration graph is shown in Fig.˜2.
@@ III-A Detect Unsafe & Unreliable Behaviors
Algorithm˜1 shows the generic arbitration algorithm from [undef] that we have extended with a verification logic. The verifier is domain-specific and may run various checks on the behavior command. In Pac-Man for example, the verifier could check whether the command would lead to a collision with walls or ghosts. In autonomous driving, one could verify if the behavior output is free of collisions and respects traffic rules. Similarly, a robotic manipulator could verify if the command is within the workspace of the robot and does not exceed joint limits. More generally, the verifier could check if the format of the command is correct and respects the specification of the system. Integrating the verification step into the arbitration algorithm allows for the use of different verifiers at various levels within the arbitration graph. This enables a more fine-grained and adaptable safety concept. For example, verifiers can be customized for specific scenarios, or computationally intensive verifiers can be reserved for higher levels of the arbitration graph.
The verification step is embedded as follows: First, in Algorithm˜1, the current situation is determined. Based on this, the root arbitrator determines its best applicable and safe action using the BestOption() function. For this, it determines the set of options which are applicable in the present situation . Like in the original arbitration algorithm, the options of an arbitrator can themselves be arbitrators or behavior components. The applicable options are then sorted into a descending list according to the underlying strategy. Now, for each option , it is checked whether its command withstands a verification . If so, this option is returned as the best applicable and safe option. If none of the options passes the verification step, the arbitrator returns the error value NO_SAFE_OPTION. Otherwise, the command returned by the root arbitrator can be executed and considered safe given the assumptions of the verifier .
@@ III-C Example: Safe Pac-Man
While the original arbitration graph shown in Fig.˜2 might have been sufficient under normal conditions, it might fail in practice if a behavior component returns an unsafe or unreliable command. Consider the scenario depicted in Fig.˜3. Since the ghosts are relatively far away and there is only one dot cluster left, the only applicable behavior is EatClosestDot. Ideally, this behavior has the intended effect and Pac-Man should move towards the last remaining dots. In our hypothetical scenario, however, a bug in the underlying path planning algorithm leads to erroneous results. Consequently, the behavior component fails to generate a valid command, causing the system to behave unpredictably or even crash.
Fig.˜4 shows the arbitration graph from Fig.˜2 extended with fallback layers. With EatClosestDot not returning a valid command, the system falls back to the newly added MoveRandomly behavior component. Moving Pac-Man in a random direction is a simple action which might help the primary behavior components to escape a deadlock and find a new valid command. In our scenario, however, MoveRandomly returns a command which would lead to a collision with a wall. Therefore, this command is rejected by the verifier as well, and the system falls back to the last resort behavior StayInPlace. While it might be impossible to complete the level with this behavior, the system remains in a predictable and safe state giving the primary behavior components a chance to recover.
@@ IV Safe and Reliable Behavior Arbitration for Automated Vehicles
Arbitration graphs can be applied to a wide range of real-world scenarios. In [undefa], they have been used in the context of automated driving. There, a behavior component represents driving maneuvers of approximately to seconds such as lane changes, parking maneuvers, or crossing intersections.
@@
This section summarizes our work in [undefp], which extends [undefa] by introducing a safety concept for arbitration graphs in the context of automated driving.
@@ IV-A Arbitration Graph — What to do?
Fig.˜5 depicts a simplified arbitration graph for automated driving that will be used to showcase our safety concept. It contains three behavior components for basic lateral behaviors FollowLane, ChangeLaneLeft and ChangeLaneRight, a parking behavior ParkNearGoal and a stopping behavior EmergencyStop. The arbitration graph for a real-world application would include more behavior components and arbitrators in order to cover a wide range of driving maneuvers and scenarios. Refer to [undefq] for a detailed discussion about behavior competencies and operational design domains.
@@ IV-B Behavior Components — How to do it?
In our implementation, each behavior component has access to an environment model (situation ) and returns a trajectory (command ). The environment model contains information such as the ego vehicle state, other traffic participants, and the planned route. Fig.˜6 illustrates an example scenario with the ego vehicle in blue, another vehicle in green, and the planned ego trajectory in dashed blue.
@@
See [undefp] for more details and other behavior components for automated driving.
@@ IV-C Verification — What is considered safe?
In order to achieve vehicle safety, we provide worst-case occupancy predictions of other traffic participants and extend the behavior command with a fail-safe trajectory [undefr, undefs]. The fail-safe trajectory is designed to be safe under worst-case assumptions and can be used in fallback layers in case of a failure or too risky situation. As a result, each behavior component is responsible for generating both a planned trajectory and a fail-safe trajectory (see Fig.˜6). The safety verifier checks if the occupancy of the fail-safe trajectory does not overlap with the worst-case occupancies of other traffic participants, i.e., the fail-safe trajectory is collision-free under worst-case assumptions.
@@ IV-D Fallback Layers — How to always drive safe?
In order to mitigate unsafe or unreliable behaviors, we extended the arbitration graph in Fig.˜5 by the proposed verifiers and fallback layers highlighted by underlined text.
@@ IV-E Experiments
We validate the proposed concept in the automated driving simulator CoInCar-Sim [undeft] using handcrafted but realistic driving scenarios from our real-world test track in Karlsruhe, Germany. We analyze two different use cases in [undefp]:
@@
Here, we focus on the latter, more critical case. Traditional decision-making approaches select maneuver options solely based on their preconditions. As a result, they might select unsafe maneuver options, if the preconditions have been designed too optimistically. Fig.˜7 shows such a scenario, where the ego vehicle wants to change lanes, but another vehicle follows too closely in the target lane. We compare the evolution of the scene with and without verification in the decision-making process.
@@ IV-F Results
To provoke a risky situation, the invocation condition of ChangeLaneLeft was made overly optimistic. As a result, the arbitration graph without verification selects ChangeLaneLeft at leading to a collision at (Fig.˜7).
With verification, the safety verifier detects the risk of a collision based on the worst-case occupancies of the other vehicle (Fig.˜8). Consequently, ChangeLaneLeft fails the verification and the UrbanDriving arbitrator selects FollowLane as its next best option (Fig.˜9). This makes the ego vehicle slow down and not change the lane until the green vehicle has passed (Fig.˜10). A collision has been prevented while maintaining a smooth driving behavior.
@@ V Conclusion
This paper has presented an extension to the arbitration graph framework that focuses on improving the safety and robustness of autonomous systems operating in complex, dynamic environments. It builds upon the strengths of arbitration graphs, which provide a flexible, scalable, and transparent decision-making framework for autonomous systems. By embedding a verification step into the arbitrators and adding structured fallback layers to the arbitration graph, the proposed method ensures that only verified and safe commands are executed.
The introduced method has been demonstrated using a Pac-Man simulation, where the arbitration graph successfully maintained safe operation of the agent even in the presence of unexpected faults or bugs. Further validation has been conducted in the context of autonomous driving, where the method has demonstrated a reduction in accident risk and an improvement in system safety. These results underline the applicability of the proposed approach to real-world problems, confirming that arbitration graphs, when equipped with safety mechanisms, can effectively manage the complexities and uncertainties of autonomous decision-making.
@@
By explicitly defining the conditions under which a behavior component is considered safe, the responsibility for system safety is shifted to the verifiers used by the algorithm. This marks a crucial advancement towards safe autonomous systems, as the overall safety of the system now primarily depends on the assumptions made by these verifiers.